Using linear combinations, how would you solve:
m+3n=2
-m+2n=3

Answer by jim_thompson5910(35256)

Add the equations to get

m+3n=2
-m+2n=3
--------
0m+5n=5

5n = 5

n = 5/5

n = 1

Now use this to find m

m+3n=2

m+3(1)=2

m + 3 = 2

m = 2 - 3

m = -1

So the solutions are m = -1 and n = 1